If r varies inversely as t, and r=132 when t=4, find r when t=32.
Answer by Cromlix(4381)

Hi there,
If r varies inversely as t
r = k/t
Where k = constant of proportionality
132 = k/4
k = 4 x 132
k = 528
Formula
r = 528/t
Using formula
r = 528/32
r = 16.5
Hope this helps :-)